OpenStudy (anonymous):

y = 1 − 3^(2x − 2) what is the intercepts and domain and range

OpenStudy (mertsj):

\[y=1-3^{2x-2}\]

OpenStudy (mertsj):

Is that the equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok well to find y intercept you sub in 0 as x

OpenStudy (anonymous):

then

OpenStudy (mertsj):

y intercept is 8/9

OpenStudy (mertsj):

\[-1=-3^{2x-2}\] \[1=3^{2x-2}\] \[\log(1)=(2x-2)\log(3)\] \[0=.477(2x-2)\] \[0=2x-2\] x=1

OpenStudy (anonymous):

YES! I was just searching some log properties

OpenStudy (anonymous):

the domain should be x e R

OpenStudy (mertsj):

Domain is all reals. 3^(2x-2) will always be positive. 1- a positive number will always be less than 1 so the range is reals <1